Circular No. 6721

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970815 T. E. Harrison, B. J. McNamara, J. J. Johnson, and P. A. Mason,

New Mexico State University; D. D. Balam, University of Victoria; and A. R. Klemola, Lick Observatory, report: “We have used the infrared camera (GRIM II) on the Astrophysical Research Consortium 3.5-m telescope to image the central 6′ × 6′ of the RXTE ASM error box for GRB 970815 (IAUC 6718), and the 1.82-m Plaskett telescope of the National Research Council of Canada to obtain deep R-band (limiting R = 23) images of a 10′ × 10′ field centered on the ASM burst position. The first J-band image mosaic was obtained on Aug. 16.27 UT. The second mosaic was made on Aug. 17.39. In the 24 hr between mosaics, no sources showed significant variability. R-band images were obtained on Aug. 16.21, 16.32, 18.25, and 19.22. No new sources or highly-variable objects were identified. Comparison of the J- and R-band images reveals two objects with interesting R − J colors. The first of these is a blue object at α = 16^{h}10^{m}03^{s}.6, δ = +81^{o}32′35′′.3 (equinox 2000.0), and the second is a red source located at position end figures 00^{s}.5, 42′′.5. Neither source exhibited measurable variability. Additionally, we have identified a peculiar POSS object at position end figures 21^{s}.0, 31′43′′.0. This object is visible on the POSS E-775 plate, but not on the E-1363 plate or on either of the O plates; it is also not present in our J or R images. The peculiar POSS object does not appear to be a plate flaw.”

K. Z. Stanek, D. D. Sasselov, and M. R. Garcia, Harvard-Smithsonian Center for Astrophysics; and C. R. Robinson, Universities Space Research Association and Marshall Space Flight Center, communicate: “Analysis of images centered on the error box of the x-ray counterpart to GRB 970815 (IAUC 6718) show no significantly variable objects to a magnitude limit of R = 21. The variability for the 200 point sources in the field with 15.5 < R < 21 is rms < 0.2 mag. The images were obtained with the 1.2-m telescope of the F. L. Whipple Observatory on Aug. 16.31, 17.16, 17.22, 19.12, 19.19, 19.31, and 20.12 UT in the Cousins R filter. The images cover a 11′ × 11′ area and are available as FITS files by anonymous

ftp at cfa-ftp.harvard.edu, in pub/kstanek/GRB970815, or through the
WWW at http://cfa-www.harvard.edu/~kstanek/. Visual inspection

of the images obtained by D. D. Balam (see item above) confirms this conclusion. These images are available as FITS files by anonymous ftp at

astroftp.phys.uvic.ca, in /pub/balam/RXTE6335.”

1997 August 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green